### v2

# arXiv:2508.19585v2[econ.TH] 02 Mar 2026

Abstract
Decision makers sometimes cannot observe the consequences of their actions ex-post. This paper axiomatically characterizes a decision model in which the decision maker cares about verifying that a good consequence has been achieved. Preferences over acts identify a set of events the decision maker expects to verify. Decision makers choose acts maximizing, in expectation over verifiable events, the worst-case utility consistent with each event. A dual model captures decision makers who instead seek to obscure poor outcomes from verification. As an application, firms choosing carbon-reduction technologies may prefer less efficient but more verifiable technologies to prove emission reductions to stakeholders.
Keywords: Verifiability, uncertainty, Choquet expected utility, greenwashing, principal-agent
JEL Classification: D81, Q54, M14
1 Introduction
In many real-world decisions, the final consequences are only partially observable. For instance, when a firm chooses a method for CO2 emission reduction, the actual amount of CO2 removed from the atmosphere is not directly observed but must be inferred from indirect evidence, such as carbon compensation certificates. This lingering ex-post uncertainty about consequences can motivate choices that deviate from standard expected utility theory. A firm might not simply aim to maximize the expected reduction in emissions; its incentives are often more complex due to the presence of stakeholders. The firm may want to prove that a significant emission reduction has been achieved, or conversely, it may want to avert blame if stakeholders can prove a desirable target was missed.
This paper’s central theoretical contribution is to model these motivations by providing an axiomatic foundation for preferences that depend on the ex-post verifiability of outcomes. We ask: how does the anticipation of what can be proven after the fact shape choice under uncertainty? We show that behavior driven by a preference for or against ex-post verifiability can be naturally understood through the lens of ambiguity aversion models fulfilling comonotonic independence [undefap], which have traditionally been motivated by ex-ante uncertainty. We rationalize ambiguity attitudes as a coherent response to unobservable consequences.
The applications we consider, such as corporate environmental claims, are characterized by a lack of objective probabilities. This subjective environment often gives rise to phenomena like “greenwashing,” where claims about expected emission reductions greatly differ between firms and stakeholders. Consequently, we adopt a fully subjective Savage framework with subjective mixtures [undefae] to identify risk preferences.
We propose two models that capture the principal-agent relationships that can arise from imperfect ex-post verifiability using a dual-self representation similar to [undefy]. The first model, certification utility, describes a decision-maker who wants to prove a good outcome was achieved. They evaluate acts based on the expected utility of the outcome they can prove, i.e., the worst outcome consistent with their best proof. This captures the behavior of a firm genuinely seeking to demonstrate its positive environmental impact and corresponds to extreme ambiguity aversion. The second model, obfuscation utility, captures the opposite motive. Here, a stakeholder presents evidence, and the firm tries to point to the most favorable interpretation consistent with that evidence. This model provides a behavioral foundation for greenwashing, where firms exploit ambiguity to avoid blame for poor outcomes, and corresponds to extreme ambiguity seeking.
Our main representation theorems show that from the decision maker’s choices alone, an outside observer can identify the key elements of the model: the presence of a hidden stakeholder, the nature of their ex-post interaction (verification or obfuscation), and the set of events the decision maker anticipates being verifiable along with the probabilities of these events. This means we can distinguish between a firm that is an expected utility maximizer, one that is genuinely seeking verifiability, and one that is engaging in obfuscation, simply by observing their preferences. The endogenously derived set of verifiable events aligns with standard assumptions in contract theory, being closed under intersections but not necessarily under complements.
Our decision-theoretic results have policy implications for corporate carbon responsibility (CCR). As policymakers seek to regulate unverified environmental claims—for instance, by planning to “proscribe […] generic environmental claims […] without proof” and “claims based on emissions offsetting schemes” [undefac]—our framework provides a tool to analyze the behavioral consequences. A welfare analysis reveals that both certification-seeking and obfuscating behaviors can lead to welfare losses. These losses are non-monotonic in the degree of verifiability; that is, increasing transparency requirements may lower welfare unless full ex-post verifiability is achieved. This suggests that policies promoting CCR and transparency must be designed with care, as their effectiveness depends critically on the set of actions available to firms.
The paper proceeds as follows: Section 2 introduces the example application of carbon offsetting programs in more detail. Section 3 presents the notation before Section 4 discusses the two decision models. The axioms that characterize these models are developed in Section 5. Comparative statics for the models can be found in Section 6. The paper also contributes a decision-theoretic foundation to a model of greenwashing and on verifiability in contract theory. Such relations to the literature are discussed in Section 7.
2 Introductory Example
This section introduces an example of how the model presented in this paper can capture certain stylized facts [^1] In the U.N. Race to Zero Campaign [undefar], an analysis of the largest firms of eight sectors [undefaj] found that most firms relied on renewable energy certificates to offset carbon emissions while only a small number of firms used certificates of nature-based carbon capture. [undefaj] concluded that “Companies’ climate pledges for 2030 fall well short of the required ambition and are inappropriately verified” (p. 6), that “the climate strategies of 15 of the 24 companies [are] of low or very low integrity” and that “targets and potential offsetting plans remain ambiguous” (p. 5). about how firms interact with their stakeholders in CCR issues where beliefs are subjective and consequences are imperfectly observable, such as carbon emission reductions. In such settings, subjective beliefs (for example about the efficacy of the firm’s actions) may greatly differ between stakeholders and firms and there usually does not exist an objective probability distribution to relate to. If firms and stakeholders cannot agree ex-ante what the best course of action is, this leaves only the ex-post outcomes for the stakeholder to evaluate whether the firm has done well.
Suppose a firm decides between different carbon offsetting programs. One option is to purchase certificates of nature-based carbon capture, or more pointedly, to plant Trees. Another option is to purchase renewable energy certificates, RECs, and the third option is to improve the production processes to increase Efficiency. For simplicity, we assume that the total budget to be spent on the offsetting programs is fixed.
Suppose the set of consequences $\mathscr{X}=\mathbb{R}_{+}$ is the amount of CO2 emissions avoided in megatons. Let the set of states $\mathscr{S}=\{s,t,u\}$ consist of three mutually exclusive states. In state $u$, there is a low supply of emission-reduction certificates (both RECs and nature-based carbon removal). This makes purchasing such certificates costly and the budget is only sufficient to offset a small amount of CO2 emissions. In states $s$ and $t$, the supply of such certificates is sufficient to offset a large amount. In state $s$, purchasing RECs leads consumers in the energy market to substitute from consuming a mix of sustainably and unsustainably produced electricity to a higher share of unsustainably produced energy. In state $t$, no such shift in behavior occurs. The options are shown in Figure 1.
Act $s$ $t$ $u$
Trees 70 70 10
RECs 60 100 10
Efficiency 40 40 40
Figure 1: Carbon Reduction Programs
Which of the three programs should a firm choose? According to expected utility, this would depend on the risk preferences and the subjective beliefs of the decision maker. A sufficiently high subjective probability of $s$, $t$, or $u$ can induce any of the three acts Trees, RECs, or Efficiency, respectively, to be optimal.
Suppose for a moment that there is no way of observing (or proving) the exact amount of CO2 reduction achieved by the different acts because only the total amount of emissions from many polluters can be measured in the atmosphere. In the absence of objective probabilities over states and when consequences are completely unverifiable, we can think of two stereotypical behaviors of firms when interacting with their stakeholders about carbon emissions:
A greenwashing firm may choose to purchase RECs and may exaggerate in their communication to stakeholders the likelihood that state $t$ obtains. For example, [undefaj] claim that for most firms joining the U.N. race to zero campaign [undefar], the CO2 emission reductions are not sufficiently verified. Firms may choose to purchase RECs even if they subjectively believe state $t$ to be very unlikely. Such behavior can be detected by an analyst from preferences under which uncertain acts are indifferent to their best possible consequence.
A certification-seeking firm may choose Efficiency to eliminate any ex-post doubt of their stakeholders that they have achieved a certain amount of CO2 emission reductions. For example, a firm may try to prove to its stakeholders that its products are carbon neutral. Such firms may choose Efficiency even if they believe state $u$ to be very unlikely because ex post they cannot exclude this state. If however the event $\{s,t\}$ would be ex-post verifiable, they may choose Trees instead. Such behavior can be detected from preferences that treat uncertain acts indifferent to their worst possible consequence.
In the following, we provide two models —and axiomatic characterizations thereof— that correspond to these two types of behavior. Compared to the extreme cases discussed above, we allow for some events, for example $\{s,t\}$, to be ex-post verifiable and subjectively determined. From an applied perspective, the main contribution is to provide conditions on preferences over emission reduction methods from which greenwashing behavior can be distinguished from certification-seeking or expected utility behavior. In addition, for any firm following either preference, we can identify what information they expect to become available in the future.
3 Notation
Let $\mathscr{X}$ be a set of consequences and $\mathscr{S}$ a finite set of states of the world. $\mathscr{E}\subseteq 2^{\mathscr{S}}$ denotes the sigma algebra of events. A simple act is a measurable function $a:\mathscr{S}\rightarrow\mathscr{X}$ with a finite image. If $E$ is an event and $a,b$ are acts, then $a_{E}b$ is the act that agrees with $a$ on all states $s\in E$ and that agrees with $b$ on all states $s\in\overline{E}\equiv\mathscr{S}-E$. If $x\in\mathscr{X}$, then $x$ also denotes the constant act that implements $x$ in all states.
[…]
An event $E$ is said to be null if $\beta\sim\gamma E\beta$ for all $\gamma,\beta\in\mathscr{X}$ such that $\gamma\succ\beta$. An event $E$ is said to be universal if $\gamma\sim\gamma E\beta$ for all $\gamma,\beta\in\mathscr{X}$ such that $\gamma\succ\beta$. An event $E$ is said to be essential if $\gamma\succ\gamma E\beta\succ\beta$ for some $\gamma,\beta\in\mathscr{X}$.
Unlike in expected utility, in our axiomatization what happens on null events may in principle still be preference relevant. We therefore introduce the stronger notion of irrelevant events. An event $E$ is said to be irrelevant if $\gamma_{E}a\sim a$ for all $\gamma\in\mathscr{X}$ and $a\in\mathscr{A}$. A state of the world is said to be irrelevant if it is an element of an irrelevant event. The set of events that are not irrelevant is denoted $\mathscr{E}^{*}$ and the set of states that are not irrelevant is denoted $\mathscr{S}^{*}$. An event that is null need not be irrelevant. In our model, an event is only irrelevant for the decision maker’s preference if it is irrelevant in this stronger sense we defined here.
If $\mathscr{F}$ is a set of events then $f:\mathscr{F}\rightarrow\mathbb{R}$ is called a set function. A set function is grounded if $f(\emptyset)=0$. A set function is normalized if $f(\mathscr{S})=1$. A set function that is both normalized and grounded is called a capacity. A set function is additive if for all $E,F\in\mathscr{F}$ such that $E\cap F=\emptyset$, $f(E)+f(F)=f(E\cup F)$. A probability measure is a grounded, normalized, additive set function. A set function is called supermodular (submodular) on a set of events $\mathscr{G}$ if for all $E,F\in\mathscr{G}$, if $\eta(E)+\eta(F)\leq(\geq)\eta(E\cup F)+\eta(E\cap F)$. A set function is modular if it is both supermodular and submodular.
4 Decision Models
We analyze two decision models. In the first decision model, the decision maker wants to verify that a particular utility level has been reached. In the context of our example, this may arise because the decision maker wants to prove to a stakeholder that a certain benefit of taking the action has been achieved.
Definition 1 (Expected Certification Utility).
[…]
$U(a)=$ $\int_{s\in\mathscr{S}^{*}}\max_{E\in\mathscr{V}:s\in E}\min_{t\in E}u(a(t))d\mu$ (1)
[…]
$\mathscr{V}$ is called the set of verifiable events. If $\mathscr{V}=\mathscr{E}$, then every state is verifiable and the decision maker maximizes expected utility. The verifiable events do not necessarily form a partition but are closed under intersections and thus form a $\pi$-system that contains $\mathscr{S}^{*}$. The interpretation of a verifiable event $E$ is that if the event obtains (i.e., if the true state of the world is within $E$), then the decision maker receives a proof that they can use to prove that this event obtained. They then use this proof to convince the stakeholder that at least the consequence $\min_{t\in E}u(a(t))$ has been achieved. [^2] The model assumes the verifiable events to be act-independent. The act-dependent case is not treated in the current paper for two reasons: first, if such verifiable events would be objectively given for every act, then this would make the analysis almost trivial. Second, if there are subjectively determined act-dependent verifiable events, then almost any behavior is possible and the verifiable events are unlikely to have interesting uniqueness properties.
The decision model is a special case of an ambiguity aversion model within the intersection of Choquet expected utility and maxmin expected utility. However, the motivation for deviating from expected utility is not ambiguity aversion but the lack of ex-post verifiability of events: across verifiable events, the decision maker follows expected utility but whenever consequences arise on events that are not verifiable, the decision maker is extremely pessimistic.
Example.
Firm A would like to be able to prove to its stakeholders that a high level of reduction of CO2 emissions has been achieved. The firm believes that states $s$ and $u$ are unlikely, though not impossible. In case the event $\{s,t\}$ obtains, the firm believes it is able to prove this to its stakeholders. For example, after purchasing any amount of certificates of any kind, the firm can show these certificates to its stakeholders. However, the firm is unable to prove that $\{u\}$ obtains if certificate prices are not publicly observable – it cannot use a low number of purchased certificates to prove that it is impossible to buy more certificates. The firm is also unable to prove that $\{s\}$ or that $\{t\}$ obtains since proving this would require observing the hypothetical behavior of how other market participants would behave given a purchase of RECs or no purchase of RECs. The firm can however verify that $\{s,t,u\}$ obtains since there is no other nonnull state. Firm A therefore considers two events when making a decision, $\{s,t\}$ and $\{s,t,u\}$. With probability $\mu(\{u\})$ it can only prove that $\{s,t,u\}$ obtains and that the emission reduction is at least equal to the worst consequence of an act. With probability $\mu(\{s,t\})$ the firm can verify that $\{s,t\}$ obtains and that the emission reduction is at least equal to the worst possible consequence on states $s$ and $t$. The firm therefore multiplies the probability of $u$ with the utility of the worst possible consequence on $\{s,t,u\}$ and the probability of $\{s,t\}$ with the worst possible consequence on $\{s,t\}$. If the probability of state $u$ is sufficiently low, the firm will choose Trees. If $u$ is sufficiently likely, the firm will choose Efficiency. ∎
In the second decision model, the decision maker has the desire to obfuscate whether a bad consequence might have resulted. In the context of our example, this may arise because a stakeholder may confront the firm with evidence about the state of the world (e.g., that $\{s\}$ obtains after having chosen RECs).
Definition 2 (Expected Obfuscation Utility).
[…]
$U(a)=\int_{s\in\mathscr{S}^{*}}\min_{E\in\mathscr{V}:s\in E}\max_{t\in E}u(a(t))d\mu$ (2)
[…]
The interpretation of this model is that after a stakeholder confronts the decision maker with the proof that an event $E$ obtains, the decision maker points to the best possible consequence that might have been achieved on $E$. We assume that the stakeholder has adversarial preferences toward the decision maker, i.e., the stakeholder tries to prove that at most a certain amount of utility has been obtained. The stakeholder always uses all available verifiable information (chooses the smallest event in $\mathscr{V}$ that contains the true state) and the decision maker then always points to the best consequence on this event. This model is dual to the certification utility in the sense that for a certification utility minimizing decision maker there exists an expected obfuscation utility with a reverse preference over outcomes and vice versa.
Example.
Firm F wants to evade negative publicity about its CO2 emissions. It worries that after making their choice, some stakeholder confronts them with evidence which state of the world obtains and proving to them that a bad consequence has resulted from their chosen act. Firm F wants to make sure that given the evidence they might be confronted with, there is a state of the world consistent with this evidence in which their chosen action yields a good result. The firm expects that only state $\{s,t\}$ can be proven by the stakeholder (for example by low market prices for RECs.) However, the firm does not expect anyone to be able to obtain definitive proof that either $\{s\}$ or $\{t\}$ obtains. If $\{s,t\}$ is sufficiently likely, the firm chooses RECs in order to be able to point to the good state $t$ in which 100 megatons of CO2 have been reduced. However, if state $\{u\}$ is sufficiently likely, the firm will choose to improve its Efficiency. ∎
Both of these decision models are extreme cases and unobservable consequences may generate a rich variety of decision models that are less extreme. The present decision models only represent a starting point for the exploration of decision theories in which consequences cannot be observed. The following axioms are therefore to be understood as a starting point for a systematic analysis of preferences under uncertainty with imperfect ex-post verifiability of consequences.
5 Axioms
This section provides a set of axioms that characterize expected verification and obfuscation utilities. The axioms allow us to distinguish between verification and obfuscation seeking behavior and should be normatively compelling if consequences are not observable.
We assume the existence of a biseparable utility representation. Biseparable preferences allow us to identify the decision weights of individuals on events while making minimal assumptions about behavior. The axioms for biseparable utility are provided in [undefad] and for convenience restated in appendix I.
Axiom 0 (Biseparable Preference).
[…]
$U(\gamma F\beta)=\mu(F)U(\gamma)+(1-\mu(F))U(\beta)$ (3)
The image $U(\mathscr{X})$ is a convex set.
Thus, in the biseparable model, decision makers’ preferences are only meaningfully restricted on binary acts. Under the assumption that certainty equivalents exist and preferences are biseparable, we can define preference averages of consequences and acts. These preference averages, first introduced in [undefae], allow us to use [undefw] style axioms without resorting to objective probabilities. While objective probability lotteries can easily be implemented in a laboratory setting, for applications in the context of greenwashing behavior, such lotteries are often unavailable.
Definition 3 (Preference Average).
For all $x,y\in\mathscr{X}$ with $x\succsim y$, $z$ is a preference average of $x$ and $y$ if for some essential event $E$, $xEy\sim[{xEz}]E[{zEy}]$. $z$ is denoted by $1/2x\oplus 1/2y$. For two acts $a,b\in\mathscr{A}$, $c$ is a pointwise preference average of $a$ and $b$ if for all $s\in\mathscr{S}$, $(s)=1/2a(s)\oplus 1/2b(s). cc$ is denoted by $1/2a\oplus 1/2b$.
Example.
Suppose Trees and RECs refer to investing a fixed amount of money into the relevant carbon offset certificates. Suppose further that there are no returns to scale to either technology. If the decision maker is risk-neutral (in terms of the curvature of $u$) over the amount of CO2 emissions reduced, then $1/2$ Trees $\oplus 1/2$ RECs refers to the act in which the carbon emissions in every state of the world are the arithmetic average of Trees and RECs. This can be thought of as investing half the money into each technology. However, if the decision maker is not risk-neutral, this is not the case. Instead, in each state $1/2$ Trees $\oplus 1/2$ RECs would offer the certainty equivalent of a (perfectly verifiable) equal probability lottery of the reduction achieved by Trees and RECs on that state. ∎
Our first axiom, comonotonic independence, is a separability condition from [undefap] that is normatively appealing in our context.
Definition 4 (Comonotonic Acts).
Acts $a,b\in\mathscr{A}$ are comonotonic if for all $s,s^{\prime}\in\mathscr{S}$,
  • $a(s)\succ a(s^{\prime})\Rightarrow b(s)\succsim b(s^{\prime})$, and
  • $b(s)\succ b(s^{\prime})\Rightarrow a(s)\succsim a(s^{\prime})$.
Thus, two acts are comonotonic if they agree on the ranking of states according to whether they achieve more desirable consequences. For instance, two acts are not comonotonic if one yields a better consequence in state $s$ than in $s^{\prime}$, while the other yields a worse consequence in $s$ than in $s^{\prime}$. Constant acts are comonotonic with all acts.
Axiom 1 (Comonotonic Independence).
[…]
$1/2a\oplus 1/2c\succsim 1/2b\oplus 1/2c.$ (4)
Comonotonic independence is normatively compelling in our setting because deviations from expected utility should only arise from a lack of verifiability. A preference reversal in (4) should only occur if mixing with act $c$ makes a good consequence verifiable for one act but not the other. This is ruled out when all three acts are comonotonic. If $a,b,c$ are comonotonic, then for any verifiable event, the state with the worst outcome is the same for all three acts. Thus, mixing with $c$ does not asymmetrically change what can be verified. Analogously, also what can be obfuscated does not change: the best possible state in each verifiable event will remain the same.
The following result, a corollary of Möbius inversion of the capacity [[, e.g.,]]chateauneuf_characterizations_1989, makes precise why comonotonic independence is the right starting point for modeling verifiability in decisions under uncertainty:
Corollary 1 (Expected Verifiability Representation).
The following statements are equivalent:
  1. 1.
[…]
  2. 2.
[…]
    $U(a)=$ $p\int_{E\in\mathscr{E}^{+}}\min_{s\in E}u(a(s))dp^{+}$
    $+(1-p)\int_{E\in\mathscr{E}^{-}}\max_{s\in E}-u(a(s))dp^{-}$ (5)
    where $\mathscr{E}^{+}$, $\mathscr{E}^{-}$ are disjoint subsets of $\mathscr{E}$, $p^{+}$ is a strictly positive probability measure over ex-post information sets $\mathscr{E}^{+}$ and $p^{-}$ is a strictly positive probability measure over $\mathscr{E}^{-}$ and $u:\mathscr{X}\rightarrow\mathbb{R}$ a convex valued function.
This representation is effectively Choquet Expected Utility [[, CEU;]]schmeidler_subjective_1989. That is, a CEU-maximizing decision maker can be interpreted as maximizing an expectation over possible ex-post information sets. On each information set, the decision maker faces either a verification or an obfuscating situation: if the event belongs to $\mathscr{E}^{+}$, the worst consequence on that event is taken as given; if it belongs to $\mathscr{E}^{-}$, the best consequence is taken as given. Thus, with probability $p$, the decision maker expects a verification situation—needing to prove that a good outcome occurred—with $p^{+}$ governing the distribution over what will be verifiable. With probability 1p1-p, they expect an obfuscating position, with $p^{-}$ governing the distribution over ex-post available verifiable events and a reversed ranking of consequences. Because of this reversal of preferences over consequences, we immediately see that comonotonic independence together with a fixed ordering of outcomes requires the decision maker to either be obfuscation-seeking or certification-seeking, but not both.
While the representation nests our two main models as special cases, it imposes no structure on which events belong to $\mathscr{E}^{+}$ and which to $\mathscr{E}^{-}$. In particular, disjoint events $E,F\in\mathscr{E}^{+}$ may have their union $E\cup F\in\mathscr{E}^{-}$: the decision maker pessimistically evaluates $E$ and $F$ individually yet optimistically evaluates any act whose outcome is only known to lie in $E\cup F$. This is difficult to reconcile with a coherent model of verifiability—what can be proven ex-post should be determined by the state that occurred, not by an independent draw from $p^{+}$ and $p^{-}$ that lives outside the state space. The axioms we impose next restore this coherence.
We therefore now impose axioms on the structure of what is verifiable ex-post. Our main identification tools will be sensitive events and reactive events, that help us identify what the decision maker expects to be ex-post verifiable: If a decision maker only cares about the good consequences that they can verify ex-post, they will often be unresponsive to improvements of acts on nonnull states if these improvements are not verifiable. Conversely, if on an event the decision maker dislikes a worsening of the consequence on any state, then the decision maker must expect that if the event occurs, they will be for sure in possession of some proof that the event (or a subevent) obtained. We thus define:
Definition 5 (Sensitive event).
Let $\gamma\succ\beta$. An event $E$ is sensitive if $\gamma_{E}\beta\succ\gamma_{E-F}\beta$ for all nonnull events $F\subset E$.
Thus, an event is sensitive if getting a worse outcome on any nonnull subevent decreases preference. For an expected utility maximizer, all nonnull events are sensitive. In certification utility, an event $E$ is sensitive if and only if the verifiable subsets of $E$ form a cover of EE—that is, every state in $E$ belongs to at least one verifiable subevent of $E$.
In obfuscation utility, when a stakeholder confronts the decision maker with evidence that event $E$ obtained, the decision maker responds by pointing to the best possible consequence within $E$. Hence, even a single good outcome anywhere in $E$ provides a defense—a strictly bad consequence can only be proven if every state in $E$ yields a bad outcome. As a result, the decision maker strictly benefits from any improvement on any nonnull subevent of $E$. We thus define:
Definition 6 (Reactive event).
Let $\gamma\succ\beta$. An event $E$ is reactive if $\beta_{E-F}\gamma\succ\beta_{E}\gamma$ for all nonnull events $F\subset E$.
A reactive event therefore requires only an improvement on any nonnull event and this increases the preference of the decision maker. Again, in expected utility all events are reactive. In our obfuscation utility, reactive events are the events $E$ for which the stakeholder has a cover of verifiable events in $\mathscr{V}$ that are all subsets of $E$.
For example, if $E-\{s\}$ is verifiable but $E$ itself is not, a good outcome in state $s$ never helps the decision maker. When the true state is $s$, any verifiable event the stakeholder can use must extend beyond $E$, so the decision maker can already point to a good outcome outside $E$. When the true state is in $E-\{s\}$, state $s$ does not appear in the evidence at all.
The decision maker can combine proofs of events. This generates a combined proof that the intersection of the two events has occurred. This is the standard assumption of verifiability in contract theory [undefx].
Axiom 2 (Proof Conjunction).
[…]
To be precise, sensitive events are events for which covers of potentially many proofs exist. However, if both $E$ and $F$ have such a cover of proofs, then we can form a cover for $E\cap F$ by intersecting every verifiable event in the cover of $E$ with every verifiable event in the cover of $F$.
Example.
Suppose the firm can verify both $\{s,t\}$ (by presenting purchase certificates, showing certificate supply was sufficient) and $\{t,u\}$ (via an academic study showing that purchasing RECs in this market does not cause energy-market demand shifts). Proof Conjunction requires that $\{s,t\}\cap\{t,u\}=\{t\}$ is also a sensitive event. The firm can present both proofs simultaneously when state $t$ is the true state: the certificate proof of $\{s,t\}$ rules out state $u$, while the study proof of $\{t,u\}$ rules out state $s$. Together, they certify that the true state is precisely $\{t\}$, where RECs achieves 100 megatons compared to 70 for Trees. If state $t$ is sufficiently likely, this additional proof therefore makes RECs the preferred choice over Trees. ∎
The final key axiom makes precise that the decision maker is an expected utility maximizer under the condition that every consequence is verifiable. We impose this via ambiguity neutrality from the literature on (ex-ante) ambiguity attitudes.
Axiom 3 (Proof Sufficiency).
[…]
This axiom states that if an event is covered by two sensitive events $E$ and $F$, then the bets on $A\cap E$, i.e., act $f$ and bets on $A\cap F$, i.e., act $g$, are ambiguity neutral, i.e., there is no motive for hedging; the mixed act $1/2f\oplus 1/2g$ is indifferent to both $f$ and $g$.
Example.
Again suppose both $E=\{s,t\}$ and $F=\{t,u\}$ are sensitive. Consider first $A=\{s,t,u\}$, so $f=\gamma_{\{s,t\}}\beta$ and $g=\gamma^{\prime}_{\{t,u\}}\beta^{\prime}$. The good outcomes of both acts fall on verifiable events: the certificate proof establishes $\gamma$ for $f$ whenever $s$ or $t$ obtains, and the study proof establishes $\gamma^{\prime}$ for $g$ whenever $t$ or $u$ obtains. The consequences of both acts are fully verifiable in every state and are therefore evaluated like expected-utility acts and mixing these acts should provide no benefit.
Now consider $A=\{s,u\}$, so $f=\gamma_{\{s\}}\beta$ and $g=\gamma^{\prime}_{\{u\}}\beta^{\prime}$: the good outcomes now fall on non-verifiable states and indifference of the two acts requires $\beta\sim\beta^{\prime}$. Under $f$, the firm achieves $\gamma$ in state $s$, but the certificate proof of $\{s,t\}$ only certifies the worst outcome on $E$, which is $\beta$ in state $t$. The good outcome is not verifiable. The same applies to $g$ in state $u$: the study proof of $\{t,u\}$ certifies at most $\beta^{\prime}$, again because state $t$ yields $\beta^{\prime}$ under $g$. State $t$ is a bottleneck in both proofs. Mixing the two acts does not resolve this: state $t$ still yields a bad outcome $1/2\beta\oplus 1/2\beta^{\prime}$. ∎
5.1 Characterization of Expected Certification Utility
The above stated axioms contain necessary and sufficient conditions on preferences for an expected certification utility. More precisely, we obtain the following characterization result:
Theorem 1 (Representation Theorem).
The following statements are equivalent:
  1. 1.
[…]
  2. 2.
[…]
The representation has standard uniqueness properties with respect to $U$ and $u$. The interesting uniqueness properties are with respect to the set of verifiable events and subjective beliefs $\mu$. Can we identify from behavior which events the decision maker expects to be able to verify? It turns out that while the set $\mathscr{V}$ is not unique, its smallest elements are unique and once we close $\mathscr{V}$ under unions, the resulting set is unique and so are the beliefs over elements of this set and complements thereof:
Proposition 1 (Uniqueness).
Suppose $\succsim^{1}$ and $\succsim^{2}$ are expected certification utilities. Then $a\succsim^{1}b\Leftrightarrow a\succsim^{2}b$ for all $a,b\in\mathscr{A}$ if and only if:
  • $U^{1}=\theta U^{2}+\phi$,
  • $u^{1}=\theta u^{2}+\phi$,
  • $cl_{\cup}(\mathscr{V}^{1})=cl_{\cup}(\mathscr{V}^{2})$,
  • $\forall E\in cl_{\cup}(\mathscr{V}^{1}):\mu^{1}(E)=\mu^{2}(E)$ and $\mu^{1}(\overline{E})=\mu^{2}(\overline{E})$.
for some $\theta\in\mathbb{R}_{+}$ and some $\phi\in\mathbb{R}$
Example.
In our example, the verifiability of $\{s,t\}$ and $\{s,t,u\}$ can be inferred from behavior of the decision maker and need not be exogenously given. Moreover, the subjective probabilities of $\{s,t\}$ and $\{u\}$ are uniquely determined from behavior. ∎
Effectively, when preferences over acts are identical then if one decision maker can prove that a certain event obtains, the other decision maker is able to prove that this event obtains or even able to prove that a strict subset of this event obtains. Thus, it could be the case that decision maker 1 is in possession of proofs that $E,F,E\cup F,E\cap F$ obtain (whenever they actually obtain) while decision maker 2 is only able to prove that $E,F,E\cap F$ obtain. In terms of behavior, we would not see a difference because the decision maker will always want to provide the most precise proof possible. Thus, we cannot identify each proof uniquely but only the collection of all proofs by closure under unions, the sensitive events. Of course, this corresponds to a unique minimal cover of the state space with verifiable events.
5.2 Characterization of Expected Obfuscation Utility
We also obtain a dual theorem that characterizes obfuscation utility. Since an expected obfuscation utility is equivalent to an expected certification utility with a reverse order of preferences, supermodularity is replaced by submodularity and proof sufficiency holds with respect to reactive instead of sensitive events.
Theorem 2 (Representation Theorem).
The following statements are equivalent:
  1. 1.
[…]
  2. 2.
[…]
Naturally, the expected obfuscation utility has the same uniqueness properties as expected certification utility. Expected obfuscation utility can be interpreted as the decision maker minimizing the utility of a stakeholder with expected certification utility who has a reverse order of preference of consequences. Similarly, expected certification utility can be seen as minimizing the utility of a stakeholder who wants to nitpick, i.e., who wants to point at the possibility that a bad consequence may have been obtained by the decision maker.
6 Model Properties
In this section we formally derive the connection of verification and obfuscation preferences to ambiguity attitudes and comparative statics of the model.
Definition 7 (Ambiguity Attitude).
A decision maker is ambiguity averse if for any two acts $f,g$ with $f\sim g$, we have $\frac{1}{2}f+\frac{1}{2}g\succsim f$. A decision maker is ambiguity seeking if for any two acts $f,g$ with $f\sim g$, we have $\frac{1}{2}f+\frac{1}{2}g\precsim f$.
We did not impose any preference for or against ambiguity in our axioms but merely that for certain binary acts the decision maker is ambiguity neutral. Nonetheless, it turns out that the axioms are sufficiently strong to guarantee ambiguity averse behavior for verification preferences and ambiguity seeking behavior for obfuscation preferences:
Corollary 2 (Ambiguity Attitude).
A decision maker with an expected verification (obfuscation) utility is ambiguity averse (seeking).
Thus, our expected certification utility model is a special case of both Choquet expected utility with a supermodular capacity [undefap] and maxmin expected utility [undefag].
For performing comparative statics, we first define comparative risk preferences in the standard way. This will help us to contrast comparative risk aversion with other comparative results.
Definition 8 (Comparative Risk Preference).
$\succsim^{2}$ is at least as risk averse as $\succsim^{1}$ if whenever $1/2x\oplus 1/2y\succsim^{1}z$, then $1/2x\oplus 1/2y\succsim^{2}z$. Two decision makers are equally risk averse if each is at least as risk averse as the other.
The following result is a standard result.
Proposition 2 (Comparative Risk Aversion).
Suppose $\succsim^{1}$ and $\succsim^{2}$ are expected verification utilities or expected obfuscation utilities. Then the following statements are equivalent:
  1. 1.
    $\succsim^{2}$ is at least as risk averse as $\succsim^{1}$.
  2. 2.
    $u^{2}=t\circ u^{1}$ for some continuous, concave function $t$.
The following result shows that risk attitude is independent of the attitude towards verifiable events. Thus, the concern for verifiability cannot be captured by risk aversion and vice versa. We can see risk aversion as an aversion to ex-ante uncertainty of perfectly verifiable consequences and preference for verifiability as an aversion to ex-post uncertainty. Moreover, we only require identical preferences over two consequences in order to compare two decision makers with respect to their verifiable events.
Proposition 3 ( Comparative Statics ).
Suppose $\succsim^{1}$ and $\succsim^{2}$ are expected verification utilities with $\gamma\succ^{1}\beta$ and $\gamma\succ^{2}\beta$ and identical null events. Then the following statements are equivalent:
  1. 1.
    $cl_{\cup}(\mathscr{V}^{1})\subseteq cl_{\cup}(\mathscr{V}^{2})$.
  2. 2.
    $\gamma E\beta\sim^{2}\gamma(E-F)\beta$ implies $\gamma E\beta\sim^{1}\gamma(E-F)\beta$.
The result shows why sensitive events are at the core of preference for verifiability. If one decision maker’s set of sensitive events is a subset of the other decision maker’s, then it is also necessarily the case that their set of verifiable events is a subset.
In the context of our example, we may care about how inefficient a decision maker’s choice is when following an expected certification utility or an expected obfuscation utility. A plausible benchmark is expected utility maximization because the agent is probabilistically sophisticated over verifiable events.
Definition 9 (Welfare Loss).
[…]
$L_{\succsim,\mu}(A)=\max_{a\in A}\int u\circ ad\mu-\max_{b\in\arg\max_{a\in A}U}\int u\circ bd\mu$ (6)
It is noteworthy that from observing expected verification preferences alone we cannot uniquely determine the welfare loss because the probabilities of non-verifiable events are not uniquely determined from behavior. If some consequences are not perfectly verifiable, we need to complement the preference information with probability information of non-verifiable events.
Given a fixed probability distribution over states it may be expected that increases in transparency generally decrease the welfare loss and align behavior closer with expected utility maximization or that it is more desirable for a decision maker to be certification-seeking than obfuscation-seeking. However, policy interventions that change the set of verifiable events or that change firms’ behavior from obfuscation-seeking to certification-seeking turn out to not necessarily reduce the welfare loss given a fixed $\mu$. The following examples show that (holding fixed the verifiable events) a decision maker acting according to expected certification utility may indeed face a larger welfare loss than a decision maker acting according to expected obfuscation utility and that increasing the set of verifiable events need not lead to a welfare improvement. This is noteworthy because this implies that calls for better verifiability of CO2 reduction and/or incentives against greenwashing need to be carefully evaluated for whether they will achieve the desired purpose.
Corollary 3 (Verification and Obfuscation Welfare Loss Comparison).
[…]
  1. 1.
    For all $\mu$ and all decision problems $A$, $L_{\succsim,\mu}(A)=L_{\succsim^{\dagger},\mu}(A)=0$.
  2. 2.
    There exist $\mu$ and decision problems $A$ and $A^{\dagger}$ such that $L_{\succsim,\mu}(A)>L_{\succsim^{\dagger},\mu}(A)$ and $L_{\succsim^{\dagger},\mu}(A^{\dagger})>L_{\succsim,\mu}(A^{\dagger})$.
Even though we may intuitively find that behavior according to expected certification utility is normatively more appealing than that of expected obfuscation utility, the expected utility loss may be larger or smaller, depending on the available actions.
Example.
Suppose in the example decision problem states $s$ and $u$ are very unlikely. In this case, an expected utility maximizer would choose RECs, the same action as a maximizer of expected obfuscation utility. A maximizer of expected certification utility would choose Trees and incur a welfare loss. This is because the decision maker fears being blamed for an ex-post suboptimal action that was ex-ante optimal. ∎
Definition 10 (Comparative Loss from Intransparency).
[…]
$T_{\succsim,\mu}(A,\mathscr{V}^{\prime})=L_{\succsim,\mu}(A)-L_{\succsim^{\prime},\mu}(A)$ (7)
[…]
The following result is obvious, but highlighted for its policy relevance:
Corollary 4 (Welfare Loss Compared to Perfect Information).
The welfare loss compared to perfect information is always nonnegative, $T_{\succsim,\mu}(A,2^{\mathscr{S}})\geq 0$.
It follows that a policy maker who could influence the set of verifiable events would always want to implement ex-post certainty about the consequences. The question is whether more generally any information increase reduces the welfare loss. Again, it is possible to find cases in which an information increase may increase the welfare loss.
Corollary 5 (Indeterminacy of Welfare Loss).
For all $cl_{\cup}\mathscr{V}\subset cl_{\cup}\mathscr{V}^{\prime}\neq 2^{\mathscr{S}}$, there exist decision problems $A$, $A^{\prime}$ such that $T_{\succsim,\mu}(A,\mathscr{V}^{\prime})<0<T_{\succsim,\mu}(A^{\prime},\mathscr{V}^{\prime})$
Example.
In our example, suppose $\mathscr{V}^{\prime}=\{\{s\},\{s,t,u\}\}$ and $\mathscr{V}=\{\{s,t,u\}\}$. Let $A$ contain only the act Trees and RECs with a slight payoff increase of RECs in state $u$. Suppose the beliefs are such that $\mu(\{t\})=.99$, meaning that an expected utility maximizer would strictly prefer RECs. Under verifiable events $\mathscr{V}$, the decision maker would prefer RECs and there is zero welfare loss. In $\mathscr{V}^{\prime}$ the decision maker is able to verify that event $\{s\}$ obtains, and now prefers Trees over RECs because in state $s$ the act Trees yields a higher payoff. However, this leads to a strictly greater welfare loss despite providing more information. Together, this and the previous example show that policy decisions regarding the incentives for transparency and efficiency of carbon emission reduction are nontrivial. While transparency and verifiability are perhaps intrinsically desirable, they may come at the cost of firms choosing less efficient CO2 reduction strategies, see [undefan] for a study on afforestation. This suggests that policies that are aimed at increasing ex-post verifiability need to be carefully examined whether the gain in transparency might be offset by a loss in efficiency — unless, ideally, the policy implements full ex-post transparency. ∎
From these ex-ante counterintuitive results it follows that policies aimed at increasing but not fully implementing transparency or inducing certification seeking behavior need to take into account the available actions of the decision makers. This result already holds without addressing limiting factors of this welfare analysis; in general, the decision maker’s preferences may not perfectly align with society’s and costs of different policies are not explicitly modeled.
7 Relation to Literature
Our model is a special case of Choquet expected utility [undefap]. While CEU has lost some of its popularity as a model of ex-ante ambiguity aversion, our interpretation as a preference for verifiability provides a clear normative motivation for it in the context of ex-post uncertainty and identifies a choice domain in which it is the natural representation. In particular, as discussed in the axioms section, comonotonic independence seems a natural starting point for other, less extreme versions of verifiability preferences than ours.
If the set of verifiable events is exogenously given, a characterization using cominimum additivity [undefak, undefal] instead of comonotonic independence is also possible. Cominimum additivity with respect to a set of events $\mathscr{V}$ requires that all acts that agree on the worst states in all of the events in $\mathscr{V}$ have to fulfill the independence axiom. Cominimum additivity provides intermediate cases between full independence and comonotone additivity.
Closely related to the representation of the present paper is the dual self model of ambiguity [undefy]. In this model, decision makers maximize an objective $U(a)=\max_{P\in\mathscr{P}}\min_{p\in P}\int u\circ adp$ where $\mathscr{P}$ is a set of sets of priors. The interpretation of the model is that there are two selves, a pessimistic one and an optimistic one. The pessimistic one evaluates the acts according to the maxmin expected utility model [undefag] by choosing the worst possible prior. However, the optimistic self determines the set of available priors that the pessimist can choose from. Our two main models are special cases of the dual self model. In a dual self interpretation of our model, the maximizer chooses a minimal cover of verifiable events. Each such cover corresponds to a set of priors via the restriction that a prior $p$ needs to fulfill $p(E)=\mu(E)$ for all events $E$ in the cover. In other words, to evaluate an act, the minimizer can choose to assign the entire probability of a verifiable event to the worst state of the verifiable event. Relative to the characterization of [undefy], our model provides conditions under which the maximization and minimization steps can be written inside the expectation.
Preference for verifiability is a form of information preference since the decision maker cares about how much information is ex post available about which consequence has been achieved. Information preferences are often modeled using a two stage approach [undefam, undefaq, undefab]. Given our decision model, the first stage would correspond to the verifiable events and the second stage to the final consequences and only the first stage is observed. If verifiable events $\mathscr{V}$ are objectively given and both ex-ante and ex-post probability distributions over consequences are objective, it is possible to perform our analysis using preferences over $\Delta\mathscr{V}\times\Delta\mathscr{X}^{\mathscr{V}}$. However, this way of modeling would severely restrict the explanatory power of our model. It does not allow to infer the verifiable events from behavior but requires the analyst to have data on the ex-post beliefs of the decision maker. In the context of our example, this would defeat the main purpose of the paper since ex-post the stakeholder and decision maker would agree on the probability distribution of the CO2 emissions.
The present paper also provides a decision theoretic foundation for the standard notion of verifiability employed in contract theory starting with [undefx]: verifiable events are closed under intersection but not necessarily under relative complements. Our application shows that this notion can be productively used to model greenwashing.
Our research also relates to definitions [undefaa] and formal models of greenwashing [undefas] and of green products [undefai]. Unlike previous modeling attempts, the present paper provides a purely behavioral definition of greenwashing versus certification-seeking behavior. We do not require unobservable model components such as incorrect consumer beliefs due to deception or detailed information about the interaction between consumers and firms. This yields a very parsimonious model of greenwashing which does not require any information about the behavior of consumers of greenwashed products – preference data of the firm over “green” policies fully identifies the model.
8 Discussion
When consequences are not directly observable to decision makers, it is plausible that these decision makers deviate from expected utility. We provide a starting point for the analysis of such deviations. As our axiomatic analysis shows, unobservable consequences can rationalize a form of ambiguity preferences. Specifically, seemingly ambiguity averse behavior can be induced from the desire to be able to prove that good consequences have been reached and seemingly ambiguity loving behavior can be induced from obfuscation seeking behavior. These models only provide extreme starting points for a more general analysis of ex-post verifiability within the CEU framework.
Our reinterpretation of CEU opens a wide field of applications for CEU when verifiability of consequences matters, such as in environmental economics. Starting with [undefar], there have been attempts to increasingly rely on CCR to achieve CO2 emission reductions. These attempts have been criticised as “inappropriately verified” [[], p.6]new_corporate_2023. Our comparative statics results suggest that —unless perfect ex-post verifiability is implemented— even in our simplistic model application the welfare effect of policies aimed at increasing verifiability of consequences are indeterminate. In such cases, traditional environmental policies (taxation of fossil fuels, energy efficiency mandates, etc.) that directly incentivize actions rather than consequences avoid this indeterminacy altogether, since their welfare effects do not depend on the structure of verifiable events.
Acknowledgements
The author gratefully acknowledges support from the Japan Society for the Promotion of Science (JSPS) under grant number 25K05000.
Appendix A Proof of Corollary 1
We prove sufficiency. By comonotonic independence and the biseparable utility representation we obtain [[, see Proposition 2 of]]ghirardato_subjective_2003 a representation $U(a)=\int(u\circ a)d\mu$ where $\mu$ is a capacity and the integral is in the sense of Choquet.
Let $m^{\mu}$ be the Möbius inverse of $\mu$, i.e., $m^{\mu}(E)=\sum_{A\subseteq E}(-1)^{|A|-1}\mu(A)$. By Proposition 2 of [undefz], the Choquet integral can be expressed in terms of the Möbius inverse [[, for a comprehensive overview, see ], p. 235]grabisch_set_2016:
$\int(u\circ a)d\mu=\sum_{E\in\mathscr{E}}m^{\mu}(E)\min_{s\in E}u(a(s)).$ (A.8)
such that $m^{\mu}(\emptyset)=0$, $\sum_{E\in\mathscr{E}}m^{\mu}(E)=1$, and all partial sums across events contained in an arbitrary event are nonnegative.
[…]
Appendix B Proof of Theorem 1
Proof.
We prove sufficiency. Via the proof of Corollary 1 we have a representation by a Choquet integral:
$\int(u\circ a)d\mu=\sum_{E\in\mathscr{E}}m^{\mu}(E)\min_{s\in E}u(a(s)).$ (B.9)
We first prove that ambiguity neutrality above sensitive events implies modularity of $\mu$ above sensitive events.
Definition 11 (Modularity above sensitive events).
A capacity $\mu$ is modular above sensitive events if for all sensitive events $E$ and $F$, and all events $A\subseteq E\cup F$:
$\mu(A\cap E\cap F)+\mu(A)=\mu(A\cap E)+\mu(A\cap F).$ (B.10)
Lemma 1 (Modularity).
[…]
Proof.
Without loss of generality, assume $u(\gamma)-u(\beta)=1=u(\gamma^{\prime})-u(\beta^{\prime})$ and $f=\gamma A\cap E\beta\sim\gamma^{\prime}A\cap F\beta^{\prime}=g$. It follows that
$U(1/2f\oplus 1/2g)=$ $1/2\mu(A\cap E\cap F)+1/2\mu(A)+\frac{u(\beta)+u(\beta^{\prime})}{2}$ (B.11)
$U(f)=$ $\mu(A\cap E)+u(\beta)$ (B.12)
$U(g)=$ $\mu(A\cap F)+u(\beta^{\prime})$ (B.13)
are all equal by proof sufficiency. But then $2\cdot U(1/2f\oplus 1/2g)=U(f)+U(g)$ or after substitution:
$\mu(A\cap E\cap F)+\mu(A)+u(\beta)+u(\beta^{\prime})=\mu(A\cap E)+u(\beta)+\mu(A\cap F)+u(\beta^{\prime})$ (B.15)
and thus $\mu$ is indeed modular above $E\cup F$.
From modularity of $\mu$ above $E\cup F$ follows that the sensitive events are not only closed under intersections, but also under unions. This is because modularity implies additivity of the Choquet integral above sensitive events $E$ and $F$: $U(1/2\gamma_{E\cap A}\beta\oplus 1/2\gamma^{\prime}_{F\cap A}\beta^{\prime})=1/2U(\gamma_{E\cap A}\beta)+1/2U(\gamma^{\prime}_{F\cap A}\beta^{\prime})$. ∎
We now prove the key lemma which relates properties of $m^{\mu}$ to whether a set is sensitive.
Lemma 2.
For all $S\subseteq\mathscr{S}^{*}$:
  • $m(S)\geq 0$, and
  • if $m(S)>0$ then $S$ is sensitive and there does not exist a cover of sensitive sets that are all strict subsets of $S$.
Proof.
We prove this by induction on the cardinality of $S$, $n$.
Case $n=1$: Since $\mu$ is a capacity, $m(S)+m(\emptyset)=\mu(S)\geq\mu(\emptyset)=m(\emptyset)=0$. Thus $m(S)\geq 0$. If $m(S)=\mu(S)>0=\mu(\emptyset)$, then since $S$ contains a single element, $S$ is sensitive and has no sensitive subsets.
Case $n>1:$ Suppose for all sets $S$ of size n1n-1 or smaller the induction hypothesis holds. We distinguish the case that $m(S)<0$ from the case that $m(S)>0$ and there exists a cover of $S$ of sensitive events that are strict subsets of $S$ and derive a contradiction for each case.
  • Suppose for sake of contradiction that $m(S)<0$.
    We first show that $S$ is sensitive and there exists a cover of sensitive subsets of $S$. For every $s\in S$ we have by monotonicity of the capacity that $\mu(S)\geq\mu(S-\{s\})$. It follows from the definition of $m$ that $m(S)+\mu(S-\{s\})+\sum_{E\subset S:s\in E}m(E)=\mu(S)\geq\mu(S-\{s\})$ and thus $m(S)+\sum_{E\subset S:s\in E}m(E)\geq 0$. Therefore for some $E\subset S$ containing $s$ we have that $m(E)>0$ and by the induction hypothesis $E$ is sensitive. It follows that every $s\in S$ is contained in a sensitive event and since by Lemma 1 sensitive events are closed under unions, $S$ is also sensitive.
    Since $n\geq 2$ and there exists a cover of sensitive subsets of $S$ and sensitive subsets are closed under unions, we can find a cover $\{A,B\}$ of two sensitive subsets of $S$ with ABA-B and BAB-A nonempty. Then by modularity above sensitive events, $\mu(S)+\mu(A\cap B)=\mu(A)+\mu(B)$. Thus, by the definition of $m$, $\sum_{E\subseteq S}m(E)+\sum_{E\subseteq A\cap B}m(E)=\sum_{E\subseteq A}m(E)+\sum_{E\subseteq B}m(E)$. This is equivalent to: $m(S)+\sum_{E\subset S:E\not\subseteq A,B}m(E)=0$. Because by assumption $m(S)<0$, this is only possible if $m(E)>0$ for some $E\subset S$, $E\not\subseteq A,B$. But from proof sufficiency follows that $\mu(E)+\mu(E\cap A\cap B)=\mu(A\cap E)+\mu(B\cap E)$ and thus $\sum_{F\subset E:F\not\subseteq A,F\not\subseteq B}m(F)<0$, contradicting the induction hypothesis.
  • Suppose for sake of contradiction that $m(S)>0$ and there exists a cover of sensitive subsets of $S$. Then $\mu(S)+\mu(A\cap B)=\mu(A)+\mu(B)$ by modularity above sensitive events. Thus, $\sum_{E\subseteq S}m(E)+\sum_{E\subseteq A\cap B}m(E)=\sum_{E\subseteq A}m(E)+\sum_{E\subseteq B}m(E)$ which is equivalent to: $m(S)+\sum_{E\subset S:E\not\subseteq A,B}m(E)=0$. But this is only possible if $m(E)<0$ for some $E\subset S$, contradicting the induction hypothesis.
Lemma 3.
If $E,F\in\mathscr{E}$, $m^{\mu}(E)>0$, $m^{\mu}(F)>0$ and $s\in E\cap F$, then there exists $G\in\mathscr{E}$ such that $G\subseteq E\cap F$, $m^{\mu}(G)>0$, and $s\in G$.
Proof.
By Lemma 2, $E$ and $F$ are sensitive. It follows from proof conjunction that $E\cap F$ is sensitive. Suppose $m(G)=0$ for all $G\subseteq E\cap F$ such that $s\in G$. Then $\mu(E\cap F)=\sum_{G\subseteq E\cap F}m^{\mu}(G)=0+\sum_{G\subseteq(E\cap F)-\{s\}}m^{\mu}(G)=\mu((E\cap F)-\{s\})$, contradicting that $E\cap F$ is sensitive. Thus, there exists some $G\subseteq E\cap F$ such that $m^{\mu}(G)>0$. ∎
Lemma 4.
For every $s\in\mathscr{S}^{*}$ there exists a unique event $\phi(s)\in\mathscr{E}$ such that $m^{\mu}(\phi(s))>0$ and if $s\in F$ and $\mu(F)>0$, then $\phi(s)\subseteq F$.
Proof.
Since $s\in\mathscr{S}^{*}$, it must be included in at least one event $E$ with $m^{\mu}(E)>0$. Since $\mathscr{S}$ and $\mathscr{E}$ are finite, there exists some $E$ such that $m^{\mu}(E)>0$ and $m^{\mu}(F)=0$ for all $F\subset E$ such that $s\in F$. To see that there exists only one such event, suppose $s\in E\cap F$, $m^{\mu}(E)>0$ and $m^{\mu}(F)>0$. Then for some $G\subseteq E\cap F$, $m^{\mu}(G)>0$. If neither $E$ nor $F$ have strict subsets on which $m^{\mu}$ is strictly positive, then by Lemma 3 it must be the case that $E=F=G$. ∎
Let $\phi:\mathscr{S}^{*}\rightarrow\mathscr{E}$ be the function that maps states $s$ into the smallest event $E\ni s$ such that $m^{\mu}(E)>0$. Let $\phi^{-1}(E)=\{t\in E|\phi(t)=E\}$ be the set of all states that $\phi$ maps into event $E$. By Lemma 4, $\phi$ and $\phi^{-1}$ are well defined.
We define a probability measure $\eta$ inductively by: $\eta(\{s\})=0$ if $s\not\in\mathscr{S}^{*}$, $\eta(\{s\})=m^{\mu}(\phi(s))/|\phi^{-1}(\phi(s))|$ if $s\in\mathscr{S}^{*}$, and $\eta(E\cup\{s\})=\eta(E)+\eta(\{s\})$ for all $E\in\mathscr{E}$ and $s\not\in E$. Thus, for every state $s$ we find the probability mass of the states $\phi^{-1}(E)$ and divide it evenly among these states.
Denote $\mathscr{V}=\{E\in\mathscr{E}|m^{\mu}(E)>0\}$, then:
$\int_{s\in\mathscr{S}}\max_{E\in\mathscr{V}:s\in E}\min_{t\in E}u(a(t))d\eta$
$\sum_{s\in\mathscr{S}^{*}}\eta(s)\max_{E\in\mathscr{V}:s\in E}\min_{t\in E}u(a(t))$
$=$ $\sum_{s\in\mathscr{S}^{*}}\eta(s)\min_{t\in\phi(s)}u(a(t))$
$=$ $\sum_{E\in\mathscr{E}}\sum_{s\in\phi^{-1}(E)}\eta(s)\min_{t\in E}u(a(t))$
$=$ $\sum_{E\in\mathscr{E}}m^{\mu}(E)\min_{t\in E}u(a(t))=U(a).$ (B.16)
The first equality sign follows from $\eta(s)=0$ for all $s\not\in\mathscr{S}^{*}$. The second equality sign follows since if $F\ni s$ and $F\in\mathscr{V}$, then by the Lemma 4, $\phi(s)\subseteq F$ and thus $\min_{t\in\phi(s)}u(a(t))\geq\min_{t\in F}u(a(t))$. The third equality sign follows since $\phi:\mathscr{S}^{*}\rightarrow\mathscr{E}$ is a well defined function and thus each state appears exactly once in the summation $\sum_{E\in\mathscr{E}}\sum_{s\in\phi^{-1}(E)}$. The fourth equality sign follows by definition of $\eta$, since $m^{\mu}(E)=\sum_{s\in\phi^{-1}(E)}\eta(s)$. ∎
Appendix C Proof of Corollary 2
Proof.
We first prove the following general lemma.
Lemma 5.
[…]
$\mu(E\cup F)+\mu(E\cap F)\geq\mu(E)+\mu(F).$ (C.17)
Proof.
Note that if an event $E$ is sensitive, removing any state from it decreases $\mu(E)$. If it is not sensitive, there exists a state $s$ such that $\mu(E)=\mu(E-\{s\})$.
If $E$ and $F$ are sensitive, supermodularity directly follows from setting $A=E\cup F$ in the definition of modularity above sensitive events.
If one of the two events is not sensitive because a state $s$ can be removed, removing such states from EFE-F and FEF-E does not change the RHS and does not increase the LHS of (C.17) because $\mu$ is a capacity. Therefore it is sufficient to prove supermodularity in the case in which all states that can be removed from non-sensitive events $E$ and/or $F$ are within $E\cap F$.
Suppose $E$ is not sensitive and we could remove state $s\in E\cap F$. Define $E^{\prime}=E-\{s\}$. $\mu(E)=\mu(E^{\prime})$. Thus, again the RHS of (C.17) stays constant if we remove $s$, i.e., $\mu(E)+\mu(F)=\mu(E^{\prime})+\mu(F)$ and the LHS may decrease. Thus, supermodularity with respect to $E$ and $F$ is implied by supermodularity with respect to $E^{\prime}$ and $F$. We can iterate this step until we reach one of two cases: Either $E$ becomes empty and supermodularity holds trivially or $E$ becomes sensitive. If $E$ becomes sensitive, we repeat the above process for event $F$. At the end of this process, either $F$ is empty and $E$ is sensitive, or both $E$ and $F$ are sensitive. In the former case, supermodularity is trivial because the LHS and RHS are identical expressions and in the latter case supermodularity follows from setting $A=E\cup F$ in the definition of modularity above sensitive events. ∎
Since an expected certification utility satisfies modularity above sensitive events (Lemma 1 in the Proof of Theorem 1) and sensitive events are closed under unions and intersections (via proof conjunction and Lemma 1), the capacity $\mu$ is supermodular.
By [undefah], a capacity is supermodular if and only if its associated Choquet integral is superadditive, i.e., for all bounded functions $h_{1},h_{2}$:
$\int(h_{1}+h_{2})\,d\mu\geq\int h_{1}\,d\mu+\int h_{2}\,d\mu.$ (C.18)
Now let $f\sim g$, i.e., $U(f)=U(g)$. The mixed act $\frac{1}{2}f+\frac{1}{2}g$ maps each state $s$ to an outcome such that $u\!\left(\frac{1}{2}f(s)\oplus\frac{1}{2}g(s)\right)=\frac{1}{2}u(f(s))+\frac{1}{2}u(g(s))$. Therefore, by superadditivity of the Choquet integral:
$U\!\left(\tfrac{1}{2}f+\tfrac{1}{2}g\right)$ $=\int\left[\tfrac{1}{2}(u\circ f)+\tfrac{1}{2}(u\circ g)\right]d\mu\geq\tfrac{1}{2}\int(u\circ f)\,d\mu+\tfrac{1}{2}\int(u\circ g)\,d\mu$ (C.19)
$=\tfrac{1}{2}U(f)+\tfrac{1}{2}U(g)=U(f),$ (C.20)
and thus $\frac{1}{2}f+\frac{1}{2}g\succsim f$.
The next lemma is trivial.
Lemma 6 (Supermodularity-Submodularity Correspondence).
[…]
By [undefah], a capacity is submodular if and only if its associated Choquet integral is subadditive, i.e., for all bounded functions $h_{1},h_{2}$:
$\int(h_{1}+h_{2})\,d\mu\leq\int h_{1}\,d\mu+\int h_{2}\,d\mu.$ (C.21)
Now let $f\sim g$, i.e., $U(f)=U(g)$. By positive homogeneity and subadditivity of the Choquet integral:
$U\!\left(\tfrac{1}{2}f+\tfrac{1}{2}g\right)$ $=\int\left[\tfrac{1}{2}(u\circ f)+\tfrac{1}{2}(u\circ g)\right]d\mu\leq\tfrac{1}{2}\int(u\circ f)\,d\mu+\tfrac{1}{2}\int(u\circ g)\,d\mu$
$=\tfrac{1}{2}U(f)+\tfrac{1}{2}U(g)=U(f),$ (C.22)
and thus $\frac{1}{2}f+\frac{1}{2}g\precsim f$. ∎
Appendix D Proof of Proposition 1
Proof.
$\Leftarrow$ is trivial, we prove $\Rightarrow$. Suppose $\succsim^{1}=\succsim^{2}$.
From the uniqueness properties of the biseparable preferences follows that $U^{1}=\theta U^{2}+\phi$, $u^{1}=\theta u^{2}+\phi$ and that the associated capacities are identical. What is left to show is the relation between the verifiable sets $\mathscr{V}^{1}$ and $\mathscr{V}^{2}$. Note that the sensitive events in the representation are exactly the sets $E$ for which $m^{\mu}(E)>0$ or for which there exists a cover of sets $E_{1},\ldots,E_{n}\subseteq E$ such that $\forall i:m^{\mu}(E_{i})>0$.
$cl_{\cup}\mathscr{V}^{1}\subseteq cl_{\cup}\mathscr{V}^{2}$: If $DV=cl_{\cup}(\mathscr{V}^{1})-cl_{\cup}(\mathscr{V}^{2})$ is nonempty, then by Lemma 1 (closure of sensitive events under unions) for every $\in DV, VV$ is sensitive in $U^{1}$. We show that $V$ cannot be sensitive in $U^{2}$: If $V$ is sensitive in $U^{2}$, then either $m^{\mu^{2}}(V)>0$ and thus $V\in\mathscr{V}^{2}$ or there exists a cover $V_{1},\ldots,V_{n}$ such that $\forall i:m^{\mu^{2}}(V_{i})>0$. If the latter is the case, then $V_{1},\ldots,V_{n}\in\mathscr{V}^{2}$ and thus $V\in cl_{\cup}(\mathscr{V}^{2})$. It follows that $V\in DV$ cannot be sensitive in $U^{2}$. But if $V$ is sensitive in $U^{1}$ but not in $U^{2}$, then the preferences cannot be identical.
$cl_{\cup}\mathscr{V}^{2}\subseteq cl_{\cup}\mathscr{V}^{1}$ follows by a symmetric argument.
$\forall E\in cl_{\cup}(\mathscr{V}^{1}):\mu^{1}(E)=\mu^{2}(E)$: Suppose for some $E\in cl_{\cup}(\mathscr{V}^{1})$, $\mu^{1}(E)>\mu^{2}(E)$. Then, for some $\gamma\succ\beta\succ\alpha$, $\gamma E\beta\succ_{1}\alpha$ and $\alpha\succ_{2}\gamma E\beta$, yielding a contradiction. ∎
Appendix E Proof of Theorem 2
Proof.
[…]
Lemma 7 (Comonotonic Independence Equivalence).
[…]
Proof.
Let $a,b,c$ be comonotonic acts. Then,
$a$ $\succsim b$ $\Leftrightarrow^{def.}$ $b$ $\succsim^{\dagger}a$
$\Leftrightarrow^{com.ind.}$ $\Leftrightarrow^{com.ind.}$
$\alpha a\oplus(1-\alpha)c$ $\succsim\alpha b\oplus(1-\alpha)c\quad$ $\Leftrightarrow^{def.}$ $\quad\alpha a\oplus(1-\alpha)c$ $\succsim^{\dagger}\alpha b\oplus(1-\alpha)c$ (E.23)
Lemma 8 (Proof Axiom Correspondence).
[…]
Proof.
[…]
$U(a)=$ $-\int_{s\in\mathscr{S}^{*}}\max_{E\in\mathscr{V}:s\in E}\min_{s\in E}u(a(s))d\mu$
$=$ $-\int_{s\in\mathscr{S}^{*}}\max_{E\in\mathscr{V}:s\in E}-\max_{s\in E}-u(a(s))d\mu$
$=$ $\int_{s\in\mathscr{S}^{*}}\min_{E\in\mathscr{V}:s\in E}\max_{s\in E}-u(a(s))d\mu$ (E.24)
Redefining the utility over consequences yields the desired result. ∎
Appendix F Proof of Proposition 2
This result is a standard result. The only interesting question is whether the definition of comparative risk preference is compatible with different verifiable events for each decision maker. Note that the definition of comparative risk preference only depends on preferences over constant acts. Further, both in expected verification utility and expected obfuscation utility the definition of the constant act $1/2x\oplus 1/2y$ does not depend on the verifiable events. Thus, risk attitudes are comparable across decision makers with different verifiable events and across decision makers with obfuscation and certification utility.
Appendix G Proof of Proposition 3
Proof.
We prove the result for expected certification utility, the derivation for expected obfuscation utility is analogous.
$1\Rightarrow 2:$ If for a nonnull $F$, $\gamma E\beta\sim^{2}\gamma(E-F)\beta,$ then $E$ is not sensitive for decision maker 2. Since the sensitive events of decision maker 2 contain all sensitive events of decision maker 1, $E$ cannot be sensitive for decision maker 1. Thus, $\gamma E\beta\sim^{1}\gamma(E-F)\beta$.
$2\Rightarrow 1:$ If $E$ is sensitive for decision maker 1 but not sensitive for decision maker 2, then for decision maker $2$ there exists some nonnull event $F\subset E$ such that $\gamma E\beta\sim^{2}\gamma(E-F)\beta$. But then also $\gamma E\beta\sim^{1}\gamma(E-F)\beta$ and thus $E$ cannot be sensitive for decision maker 1. Thus, the sensitive sets of decision maker 1 are a subset of the sensitive sets of decision maker 2.
Appendix H Proofs of Corollaries 3-5
Proof.
For the proof of Corollary 3, consider a decision problem $A$ in which the loss given a certification-seeking decision maker is nonzero. Then there exists an event $E$ such that no subset is verifiable but a superset of which is verifiable. In particular, let $F$ and $G$ be disjoint proper subsets of $E$. Consider the acts $\gamma F\beta G\delta$ and $\gamma G\beta F\delta^{\prime}$ with $\gamma\succ\beta$. Assume that the utility difference between $\delta$ and $\delta^{\prime}$ is small. An expected utility maximizing DM strictly prefers the former to the latter if and only if $\mu(F)>\mu(G)$ (up to the small utility difference between $\delta$ and $\delta^{\prime}$). A certification seeking decision maker strictly prefers the former to the latter if and only if $\beta\succ\delta\succ\delta^{\prime}$. An obfuscation seeking decision maker strictly prefers the former to the latter if and only if $\delta\succ\delta^{\prime}\succ\gamma$. Thus, it is possible that the expected utility maximizer agrees on the optimal action with the verification seeker or the obfuscation seeker without agreeing with the other.
The proof of 4 is trivial, the expected utility of the maximizer of $U$ can at most be the maximum expected utility on a decision problem. Since if every event is verifiable the decision maker maximizes expected utility, the result follows.
[…]
Appendix I Axioms for Biseparable Preferences
In the main text, a biseparable preference is assumed. Here we restate the axioms used by [undefad] to characterize biseparable preferences. The axioms are slightly adjusted to fit the notation of the present paper.
Axiom 4 (Preference Relation).
[…]
Definition 12 (Dominant Acts).
An act $f$ dominates an act $g$ if for all $s\in\mathscr{S}$, $f(s)\succsim g(s)$.
Axiom 5 (Dominance).
[…]
Axiom 6 (Eventwise Monotonicity).
[…]
  • if $E$ is nonnull, then $x\succ y$ implies $xEz\succ yEz$, and
  • if $E$ is nonuniversal, then $x\succ z$ implies $yEx\succ yEz$.
Let $\mathscr{X}$ be a topological space and let $\mathscr{X}^{\mathscr{S}}$ be endowed with the product topology.
Axiom 7 (Continuity).
[…]
Continuity ensures the existence of certainty equivalents and therefore the subjective mixtures [undefao, undefae, undefaf] are well defined:
Definition 13 (Subjective Mixture).
A subjective mixture is an act $f\oplus_{E}g:s\mapsto[{f(s)Eg(s)}]$.
Axiom 8 (Binary Comonotonic Act Independence).
[…]
$xEy\succsim x^{\prime}Ey^{\prime}\quad\Rightarrow\quad(xEy)\oplus_{B}(x^{\prime\prime}Ey^{\prime\prime})\succsim(x^{\prime}Ey^{\prime})\oplus_{B}(x^{\prime\prime}Ey^{\prime\prime}).$ (I.25)
Theorem 3 (Biseparable Utility Representation [undefad]).
[…]
  1. 1.
[…]
  2. 2.
[…]